1. A resistor R in series with an 8 ohms resistor is connected across a 60 V voltage source delivering 2.5 A of current. Find the value of resistor R.
1
Expert's answer
2022-05-10T13:21:41-0400
"R_{net}=R+8=\frac{V}{I}"

"R+8=\frac{60}{2.5}"

"R=16\;\Omega"